# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

### v1

@@ 6.2 Network Construction
Following the estimation of the CCC-GARCH, DCC-GARCH, and GO-GARCH models, we constructed six distance-based matrices to support the network-based analysis. Two raw data–based matrices were computed using Euclidean and correlation-based dissimilarities from monthly log returns. Subsequently, model-based dissimilarity matrices were computed using truncated AR($\infty$) representations of log-squared returns, following Piccolo (1990), and form the conditional correlation structures produced by the CCC, DCC, and GO-GARCH models. All matrices were constructed following the methodology detailed in the previous section, ensuring comparability and coherence in the subsequent network-modeling framework.
@@
Figure 2: Network plots of the six OPEC oil-exporting countries based on six different distance measures. Raw data-based distances: (a) Euclidean distance, (b) Correlation-based distance. Model-based distances: (c) Piccolo-based distance, (d) CCC-GARCH, (e) DCC-GARCH, and (f) GO-GARCH.
Figure 2 visualizes oil price interdependencies among the six OPEC countries through networks derived from monthly log returns. A notable contribution is the emergence of coherent and interpretable structures across all modeling approaches, which were constructed without relying on external variables. The raw-data-based networks—Euclidean (a) and correlation-based (b)—highlight regional proximity, with Algeria, Libya, and Nigeria forming a North African cluster, likely reflecting geographic and economic commonalities. In contrast, model-based networks reveal volatility-driven connections. The Piccolo-based distance (c), based on Piccolo (1990)’s AR($\infty$) structure, displays a different topology, suggesting the influence of long memory dynamics. The CCC-GARCH (d), DCC-GARCH (e), and GO-GARCH (f) networks exhibit a strong Middle Eastern core, with Saudi Arabia and the UAE emerging as central hubs, particularly Saudi Arabia, which appears as the primary source of risk spillovers to the majority of OPEC members, highlighting its clear capacity to propagate systemic risk among member countries. This dominant role aligns with its status as the largest oil producer within the organization (Fattouh, 2007; Alkhathlan et al., 2014; Al Rousan et al., 2018). Iran’s shifting position across models reflects external constraints, such as sanctions and production caps, which affect macroeconomic variables, including oil output, inflation, and government revenues (Nakhli et al., 2021). These networks, constructed solely from statistical features, highlight how spatiotemporal and GARCH-based models capture the latent interdependencies shaped by geopolitical and macroeconomic influences. These interdependencies may also reflect broader macroeconomic shocks and geopolitical frictions, as emphasized by Kilian (2008). Moreover, as noted by Mignon and Saadaoui (2024), such factors are essential for understanding systemic risk and volatility propagation within the global oil market. These visualizations identify dominant clusters and key players under varying model assumptions and serve as the structural foundation for the forecasting performance evaluation conducted in the next subsection.
@@ 6.3 Forecasting Results
First, we estimate the network-based volatility dynamics using the Generalized Method of Moments (GMM) framework proposed by Otto et al. (2024a) for dynamic spatiotemporal ARCH models. The models incorporate six different network weight matrices. Table 5 presents the estimation results for each network approach. Across all cases, we observe a pronounced instantaneous network ARCH effect ($\rho$), while the country-specific temporal ARCH effects ($\gamma_{1},\ldots,\gamma_{6}$) remain relatively small and partially statistically insignificant.
Table 5: GMM Estimation Results of the Dynamic Spatiotemporal ARCH Model under Different Network Structures
@@
Euclidean-based 0.952 -0.016 0.027 -0.015 -0.130 0.202 0.219 1.498
Std. Error 0.054 0.045 0.042 0.049 0.056 0.042 0.047
t-statistic 17.545 -0.359 0.652 -0.315 -2.320 4.827 4.623
Correlation-based 0.955 -0.014 0.021 -0.022 -0.128 0.206 0.218 1.493
Std. Error 0.054 0.045 0.042 0.049 0.056 0.042 0.047
t-statistic 17.579 -0.301 0.503 -0.442 -2.294 4.956 4.615
Piccolo-based 1.213 -0.156 0.114 -0.149 -0.024 0.342 0.388 2.541
Std. Error 0.103 0.065 0.055 0.070 0.073 0.053 0.058
t-statistic 11.791 -2.402 2.094 -2.139 -0.323 6.414 6.645
CCC-GARCH-Based 0.923 0.003 0.004 0.007 -0.070 0.175 0.121 1.587
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.050
t-statistic 17.443 0.063 0.077 0.147 -1.231 3.980 2.393
DCC-GARCH-Based 0.923 0.003 0.004 0.007 -0.070 0.174 0.122 1.587
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.050
t-statistic 17.444 0.064 0.094 0.138 -1.238 3.968 2.408
GO-GARCH-Based 0.922 0.002 0.001 0.010 -0.067 0.175 0.124 1.598
Std. Error 0.053 0.048 0.046 0.050 0.057 0.044 0.051
t-statistic 17.378 0.039 0.016 0.193 -1.184 3.980 2.454
@@
Following the estimation of the network-based volatility models, we assess their out-of-sample forecasting performance using two standard loss functions: root mean squared forecast error (RMSFE) and mean absolute forecast error (MAFE). These metrics are computed over the full forecast horizon $T$ and across all $N$ countries. Let $y_{i,t}$ denote the realized log-volatility and $\hat{y}_{i,t}$ its one-step-ahead forecast for country $i$ at time $t$. The average forecast errors are then given by:
$\mathrm{RMSFE}$ $=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{i,t}-\hat{y}_{i,t})^{2}},$ (37)
$\mathrm{MAFE}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{T}\sum_{t=1}^{T}|y_{i,t}-\hat{y}_{i,t}|.$ (38)
@@
To formally test for differences in predictive accuracy, we employ the Diebold–Mariano (DM) test (Diebold and Mariano, 1995, 2002). Let $e^{(m)}_{i,t}=y_{i,t}-\hat{y}^{(m)}_{i,t}$ denote the forecast error from model $m$. The loss differential between models $m_{1}$ and $m_{2}$ is defined as:
$d_{i,t}=L(e^{(m_{1})}_{i,t})-L(e^{(m_{2})}_{i,t}),\qquad\text{with }L(e)=e^{2}\text{ or }|e|.$
The DM test statistic is computed as:
$\mathrm{DM}_{i}=\frac{\bar{d}_{i}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d}_{i})}},\qquad\bar{d}_{i}=\frac{1}{T}\sum_{t=1}^{T}d_{i,t},$ (39)
@@
where $\widehat{\mathrm{Var}}(\bar{d}_{i})$ denotes the heteroskedasticity and autocorrelation consistent (HAC) variance estimator of Newey and West (1987), with truncation lag $k-1=0$ for one-step-ahead forecasts. Two-sided $p$-values are reported at the 5% significance level.
For comparisons involving nested models, we employ the Clark–West (CW) test (Clark and West, 2007), which adjusts for the bias introduced by overfitting in the larger model’s mean squared prediction error (MSPE). The adjusted loss differential is defined as:
$f_{t}=(y_{t+\tau}-\hat{y}_{1,t+\tau})^{2}-\left[(y_{t+\tau}-\hat{y}_{2,t+\tau})^{2}-(\hat{y}_{1,t+\tau}-\hat{y}_{2,t+\tau})^{2}\right],$
where Model 2 nests Model 1. The CW test statistic is then computed as the average of the adjusted loss differentials:
$CW=\frac{1}{P}\sum_{t=R}^{T}f_{t},$ (40)
where $P=T-R+1$ is the number of out-of-sample observations, and $R$ denotes the initial estimation window. Under the null hypothesis of equal predictive accuracy, the test statistic is asymptotically normally distributed. One-sided critical values apply: the null is rejected at the 10% significance level if $CW>1.282$, and at the 5% level if $CW>1.645$.

### v2

@@ 6.2 Network Construction
Following the estimation of the CCC-GARCH, DCC-GARCH, and GO-GARCH models using the full dataset, we constructed six distance matrices to support the network analysis. Two data-driven matrices were derived from monthly log returns using Euclidean and correlation-based dissimilarities. Subsequently, four model-based dissimilarity matrices were computed: one using the truncated AR($\infty$) representations of log-squared returns following Piccolo (1990), and three novel structures derived from the conditional correlation matrices produced by the CCC, DCC, and GO-GARCH models. All matrices were constructed following the methodology detailed in Subsection 5.2, ensuring comparability and coherence in the subsequent network-modeling framework.
@@
Figure 2: Network plots of the six OPEC oil-exporting countries based on six different distance measures. Raw data-based distances: (a) Euclidean distance, (b) Correlation-based distance. Model-based distances: (c) Piccolo-based distance, (d) CCC-GARCH, (e) DCC-GARCH, and (f) GO-GARCH. The grey colour scale of the edges is proportional to the weight degree of the connections.
The network structures depicted in Figure 2 illustrate the oil price interdependencies among the six OPEC countries. These networks were derived solely from monthly log returns. Notably, this methodology yielded coherent and interpretable structures across all modeling approaches without relying on external variables or geospatial data, underscoring the robustness of the proposed clustering approaches.
The data-driven networks—Euclidean (a) and correlation-based (b)—highlight regional proximity, with Algeria, Libya, and Nigeria forming a North African cluster, likely reflecting geographic and economic commonalities. In contrast, the model-based networks reveal volatility-driven connections. The Piccolo distance network (c), derived from Piccolo (1990)’s AR($\infty$) structure, displays a distinct topology suggesting the influence of long-memory volatility dynamics. However, this approach yields a weaker and less representative network structure, characterized by sparse connectivity and minimal edge formation. It demonstrates limited evidence of spillover effects or interconnectedness in this context. Furthermore, this topology lacks robustness, as evidenced by the inconsistent network representations across different in-sample training periods shown in Appendix Figure F.1.
Conversely, the CCC-GARCH (d), DCC-GARCH (e), and GO-GARCH (f) models provide the most accurate and consistent spatial representation of network formation compared to the heuristic measures, especially the GO-GARCH based weight matrix. These models reveal critical information about the connectivity between OPEC members, exhibiting a strong Middle Eastern core with Saudi Arabia and the UAE emerging as central hubs. Saudi Arabia, in particular, appears as the primary source of risk spillovers and the central node for the majority of OPEC members, highlighting its capacity to propagate systemic risk. This dominant role aligns with its status as the largest oil producer within the organization (Fattouh, 2007; Alkhathlan et al., 2014; Al Rousan et al., 2018). The analysis also reveals the nuanced position of Iran, whose shifting connections are influenced by external constraints—such as sanctions and production limits—that also impact its macroeconomic indicators (Nakhli et al., 2021).
These networks, constructed solely from statistical features and dissimilarity measures, demonstrate how time-series clustering methods—especially the conditional correlation MGARCH-informed networks—effectively capture latent interdependencies and volatility transmissions. These interdependencies may also reflect broader macroeconomic shocks and geopolitical frictions, as emphasized by Kilian (2008). Moreover, as noted by Mignon and Saadaoui (2024), such factors are essential for understanding systemic risk and volatility propagation within the global oil market. These visualizations identify dominant clusters and key players in the OPEC cartel, reflecting real economic dependencies under varying model assumptions. They serve as the structural foundation for the spillover channels assessed in the forecasting performance evaluation in the next subsection. (We refer the reader to Appendix Figure F.1 for a representation of the weight matrices under different initial training windows $T_{0}$ employed for the rolling window forecasting session).
@@ 6.3 Forecasting Results
First, using the in sample period of $T_{0}=450$ months (February 1983-August 2020), we estimate the network-based volatility dynamics using the Generalized Method of Moments (GMM) framework proposed by Otto et al. (2024b) for dynamic spatiotemporal ARCH models. The models incorporate six different network weight matrices estimated within the in-sample period for each model. Table 5 presents the estimation results for the in-sample for each network approach. Across all cases, we observe a pronounced instantaneous network ARCH effect ($\rho$), while the country-specific temporal ARCH effects ($\gamma_{1},\ldots,\gamma_{6}$) remain relatively small and partially statistically insignificant.
Table 5: GMM Estimation Results of the Dynamic Spatiotemporal ARCH Model (In-Sample $T=450$)
@@
Euclidean-based 0.948 -0.035 0.059 -0.053 -0.105 0.226 0.220 1.553
Std. Error 0.058 0.048 0.042 0.052 0.057 0.043 0.049
t-statistic 16.374 -0.732 1.410 -1.018 -1.840 5.223 4.528
Correlation-based 0.951 -0.031 0.050 -0.057 -0.103 0.229 0.216 1.540
Std. Error 0.058 0.048 0.042 0.052 0.057 0.043 0.048
t-statistic 16.461 -0.650 1.202 -1.090 -1.809 5.318 4.466
Piccolo-based 1.232 -0.145 0.115 -0.144 -0.066 0.344 0.362 2.503
Std. Error 0.105 0.065 0.053 0.070 0.074 0.054 0.058
t-statistic 11.726 -2.216 2.187 -2.059 -0.900 6.373 6.223
CCC-GARCH-Based 0.923 -0.033 0.044 -0.041 -0.073 0.202 0.122 1.635
Std. Error 0.056 0.051 0.045 0.054 0.058 0.045 0.051
t-statistic 16.479 -0.641 0.964 -0.757 -1.254 4.453 2.386
DCC-GARCH-Based 0.924 -0.032 0.044 -0.041 -0.073 0.201 0.123 1.635
Std. Error 0.056 0.051 0.045 0.054 0.058 0.045 0.051
t-statistic 16.479 -0.638 0.983 -0.772 -1.265 4.446 2.397
GO-GARCH-Based 0.922 -0.033 0.041 -0.039 -0.069 0.201 0.128 1.647
Std. Error 0.056 0.051 0.045 0.054 0.058 0.045 0.051
t-statistic 16.411 -0.652 0.897 -0.733 -1.182 4.412 2.484
@@
Following the estimation of the network-based volatility models, we assess their out-of-sample forecasting performance using two standard loss functions: root mean squared forecast error (RMSFE) and mean absolute forecast error (MAFE). These metrics are computed over the full forecast horizon $T$ and across all $N$ countries. Let $y_{i,t}\equiv Y_{t}^{*}(v_{i})=\ln Y_{t}^{2}(v_{i})$ denote the observed log-squared return and $\hat{y}_{i,t}$ its one-step-ahead forecast for country $i$ at time $t$. To evaluate the predictive accuracy of the full multivariate system, we compute global metrics aggregated across all countries and time periods:
$\mathrm{RMSFE}$ $=\sqrt{\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(y_{i,t}-\hat{y}_{i,t})^{2}},$ (31)
$\mathrm{MAFE}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}|y_{i,t}-\hat{y}_{i,t}|.$ (32)
To account for sampling uncertainty, we compute 95% confidence intervals for these metrics using a stationary bootstrap procedure with 2,000 replications.
@@
To formally test for differences in predictive accuracy, we employ the Diebold–Mariano (DM) test (Diebold and Mariano, 1995, 2002). Let $e^{(m)}_{i,t}=y_{i,t}-\hat{y}^{(m)}_{i,t}$ denote the forecast error from model $m$ for country $i$ at time $t$. We define the pooled loss differential between a candidate model and the benchmark (Net. GO-GARCH) across all space-time observations as:
$d_{i,t}=L(e^{(\text{bench})}_{i,t})-L(e^{(\text{cand})}_{i,t}),\qquad\text{with }L(e)=e^{2}\text{ or }|e|,$ (33)
and apply the test to the pooled sequence $\{d_{i,t}\}_{i=1,\ldots,N;\,t=1,\ldots,T}$ containing all $NT$ observations. The standard DM test statistic is computed as:
$\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\qquad\bar{d}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}d_{i,t},$ (34)
@@
where $\widehat{\mathrm{Var}}(\bar{d})$ denotes the heteroskedasticity and autocorrelation consistent (HAC) variance estimator of Newey and West (1987), with truncation lag k1=0k-1=0 for one-step-ahead forecasts.
Given the strong evidence of heteroskedasticity in our data—demonstrated by the distributional properties in Appendix Figures B.1 and B.2—and the time-varying volatility inherent in our standard MGARCH and Network Log-ARCH frameworks, the standard Diebold-Mariano test may suffer from size distortions or power loss. To address this, we employ the heteroskedasticity-robust test for equal forecast accuracy proposed by Harvey et al. (2024). When applied to our pooled panel structure, the method treats the $NT$ loss differentials as a single sequence allowing for unconditional heteroskedasticity:
$d_{i,t}=c+\sigma(i,t)\epsilon_{i,t},\quad\text{with }\mathrm{Var}(d_{i,t})=\sigma^{2}(i,t),$ (35)
where $\sigma(\cdot)$ is a deterministic positive volatility function indexed over the pooled observations and $\{\epsilon_{i,t}\}$ is a mean-zero error process. The null hypothesis of equal forecast accuracy is $H_{0}:c=0$. To test this, the method rescales the loss differential by a nonparametric estimate of the time-varying volatility, yielding the test statistic:
$\mathrm{DM}_{\text{het}}=\frac{\sqrt{NT}\cdot\overline{d^{*}}}{\sqrt{\widehat{\Omega}(d^{*})}},\quad\text{where}\quad d^{*}_{i,t}=\frac{d_{i,t}}{\hat{\sigma}_{i,t}},$ (36)
and $\widehat{\Omega}(d^{*})$ is a long-run variance estimator applied to the standardized series.
The spot volatility $\hat{\sigma}_{i,t}$ is estimated using a kernel smoother with a leave-$k$-out cross-validation scheme to handle serial correlation in the pooled sequence. Specifically, treating the $NT$ observations as indexed sequentially, the variance at each position is estimated as:
$\hat{\sigma}^{2}_{i,t}=\sum_{j}w_{j,(i,t)}d_{j}^{2},\quad w_{j,(i,t)}\propto K_{h}\left(\frac{\text{pos}(i,t)-\text{pos}(j)}{NT}\right)\cdot\mathbb{1}\{|\text{pos}(i,t)-\text{pos}(j)|\geq k\},$ (37)
where $K_{h}(\cdot)$ is a Gaussian kernel with bandwidth $h$, pos(i,t)(i,t) denotes the sequential position in the pooled series, and the indicator function ensures that the estimation excludes the $2k$ nearest neighbors. Following Harvey et al. (2024), we set $k=20$. Under the null hypothesis and standard regularity conditions, the heteroskedasticity-adjusted statistic follows a standard normal distribution asymptotically:
$\mathrm{DM}_{\text{het}}\xrightarrow{d}N(0,1).$ (38)
This adjustment ensures that the test maintains correct size even under severe volatility clustering while achieving substantial power gains over the standard DM test when heteroskedasticity is present. We report both the standard and heteroskedasticity-adjusted results, computing two-sided $p$-values at the 5% significance level.
6.3.1 Out-of-Sample Forecast Evaluation